# Disentangled Clothed Avatar Generation with Layered Representation

### v1

@@ Abstract
Clothed avatar generation has wide applications in virtual and augmented reality, filmmaking, and more. Previous methods have achieved success in generating diverse digital avatars, however, generating avatars with disentangled components (e.g., body, hair, and clothes) has long been a challenge. In this paper, we propose LayerAvatar, the first feed-forward diffusion-based method for generating component-disentangled clothed avatars. To achieve this, we first propose a layered UV feature plane representation, where components are distributed in different layers of the Gaussian-based UV feature plane with corresponding semantic labels. This representation supports high-resolution and real-time rendering, as well as expressive animation including controllable gestures and facial expressions. Based on the well-designed representation, we train a single-stage diffusion model and introduce constrain terms to address the severe occlusion problem of the innermost human body layer. Extensive experiments demonstrate the impressive performances of our method in generating disentangled clothed avatars, and we further explore its applications in component transfer. The project page is available at: https://olivia23333.github.io/LayerAvatar/
@@ 1 Introduction
Recent learning-based methods [3, 16, 69] mainly combine 3D representations [35, 58, 49, 24] with generation pipelines (e.g., 3D-aware GANs [38, 4, 5] and diffusion models [46]) to create digital avatars. However, these methods often ignore the compositional nature of digital avatars and represent the human body, hair, and clothes as a whole, which limits their capabilities in digital avatar customization such as cloth transfer. Neural-ABC [7] and SMPLicit [14] provide parametric model with disentangled clothes and human body, however, they do not model the texture. HumanLiff [21] proposes a layer-wise generation process that first generates clothed avatars in minimal clothes, then generates digital avatars wearing the next layer of clothing conditioned on the current layer. Nevertheless, the features of the human body and clothes are not fully disentangled, which makes it difficult to extract the components of each layer, thus reducing the editing ability. In addition, some methods [57, 15, 53, 60, 18] follow the trend of DreamFusion [43] to achieve disentangled clothed avatar generation by learning each component of digital avatars separately through the prior knowledge of 2D diffusion models [46] in an optimization manner. These methods can generate clothed avatars with each component disentangled, however, they take hours to generate a single digital avatar, and the training time will increase linearly according to the number of components.
@@
In this paper, we propose LayerAvatar, a feed-forward diffusion-based method that can (1) generate high-quality clothed avatars with each component disentangled, enabling transfer of clothes, hair, and shoes; and (2) generate avatars in a feed-forward manner which only requires seconds per avatar. To achieve this, we choose 3D Gaussians [24] as the underlying representation due to its efficiency in high-resolution and real-time rendering, strong representation capability, and explicit property that facilitates editing and animation capabilities. However, naively representing the disentangled clothed avatar using 3D Gaussians is impractical due to its unstructured nature that is incompatible with most current feed-forward generation pipelines [5, 46] and the difficulty in disentangling each component. Therefore, we introduce a Gaussian-based UV feature plane, in which 3D Gaussians are projected into a predefined 2D UV space shared among subjects. The attributes of each 3D Gaussian are encoded as local geometry and texture latent features, which can be obtained from the 2D feature plane via bilinear interpolation. To achieve full disentanglement of avatar components (e.g., hair, shoes, upper cloth, etc.), we represent avatar components in separate layers of the UV feature plane based on the SMPL-X [41] UV template. Given the layered representation, we can animate the generated avatar into novel poses, including gesture and facial expression controllability, and transfer each component to various body shapes by extracting prior knowledge of blend shapes and skinning weights from the SMPL-X [41] model.
@@
We evaluate LayerAvatar on multiple datasets [66, 9], demonstrating its superior performance in generating disentangled clothed avatars. The generated avatars support high-resolution (1024) and real-time rendering, and expressive pose control, including gesture and facial expression controllability. We also explore its application in component transfer. In summary, our main contributions are:
  • We introduce LayerAvatar, the first feed-forward clothed avatar generation pipeline with each component disentangled, enhancing the controllability of avatar generation.
  • We propose a layered UV feature plane representation that supports efficient rendering, expressive animation, and the disentanglement of each component.
  • Our method achieves impressive generation results on multiple datasets and facilitate the downstream application of component transfer.
@@ 2 Related work
Figure 2: Method overview. LayerAvatar learns a feed-forward diffusion model to generate clothed avatars with each component disentangled. The clothed avatars are represented as layered UV feature plane where components are represented separately. After decoding the feature plane into attribute maps, we can extract 3D Gaussians from them through SMPL-X-based templates. Generated clothed avatars are then transformed into targeted pose space for further supervision. Reconstruction loss and constrain loss are both utilized to facilitate the disentanglement and handle the severe occlusion of human body layer.
@@
Compositional Avatar Representation. Instead of representing the clothed avatar as a single entity, some methods represent clothed avatars as a combination of multiple submodules. COAP [34], Spams [40], and DANBO [52] consider human avatars as a composition of body parts, while EVA3D [19] and ENARF-GAN [39] follow this trend and utilize multiple neural networks to represent different body parts of the digital avatar, achieving more efficient and detailed generation results. Several methods [1, 64, 21] represent clothed avatars as separate layers to enable the expressiveness of various topologies. However, these methods ignore the disentanglement of human body and clothes, which makes each submodule less physically meaningful. Recently, some works [17, 42, 30, 63, 75] disentangle the human body and clothes by representing each component separately. During the rendering process, these components are combined through various compositional rendering techniques. Most of these methods are designed to optimize a single digital avatar. In this paper, we propose a novel layered UV feature plane representation that is compatible with the feed-forward generation framework.
@@ 3 Method
We propose LayerAvatar, a feedforward generative method for disentangled clothed avatar generation. The overview of our method is illustrated in Fig 2. We provide a brief introduction for prior knowledge in Sec. 3.1. To achieve disentangled clothed avatar generation, we propose a novel layered UV feature plane representation(Sec. 3.2), which facilitates disentanglement and is compatible with current feed-forward generative pipelines. The clothed avatars are generated in canonical space and then deformed to targeted pose space via deformation module (Sec. 3.3). The training process is introduced in Sec. 3.4.
@@ 3.1 Preliminary
SMPL-X [41] is an expressive parametric human model that can produce human meshes (without clothes) given shape parameter $\mathbf{\beta}$, pose parameter $\mathbf{\theta}$, and expression parameter $\mathbf{\psi}$. The producing process of human meshes $M(\mathbf{\beta},\mathbf{\theta},\mathbf{\psi})$ can be formulated as follows:
@@
Figure 3: Layered UV map and corresponding templates. We design templates to cover the maximum regions of the components.
3D Gaussians [24]. is a primitive-based explicit representation that combines the strengths of both previous explicit and implicit representations. It consists of a set of learnable 3D Gaussian primitive $\mathcal{G}_{k}$ where each contains five attributes: position $\mu$, scaling matrix $\mathbf{S}$, rotation matrix $\mathbf{R}$, opacity $\alpha$, and color $c$. In practice, we employ diagonal vector $\mathbf{s}\in\mathbb{R}^{3}$ and axis-angle $\mathbf{r}\in\mathbb{R}^{3}$ to represent $\mathbf{R}$ and $\mathbf{S}$ respectively. 3D Gaussians are represented as ellipses in 3D space defined by their position $\mu$ and covariance matrix $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$. During the rendering process, these 3D Gaussians are projected to a 2D image plane where the pixel color $C$ can be calculated as follows:
$C=\sum_{i=1}^{N}c_{i}\sigma_{i}\prod_{j=1}^{i-1}\left(1-\sigma_{j}\right),$ (2)
where $c_{i}$ is the color of the $i$-th 3D Gaussian on the ray, and $\sigma_{i}$ is the blending weight calculated with the opacity value $\alpha$ and probability density.
@@ 3.2 Layered UV Feature Plane Representation
Aiming at disentangling exterior components from human body, recent methods attempt to represent digital avatars in a compositional manner where each component is represented individually. Although these methods can disentangle clothed avatars as separate elements, they are mostly designed for optimization-based methods and are not suitable enough to adopt in a feed-forward generative scenario. Therefore, we propose a novel layered UV feature plane representation to support disentanglement of clothed avatars and is compatible with common feed-forward generation pipelines. We employ 3D Gaussians as the basic representation which displays the advantages of efficient rendering, easy animation, and editing.
@@
Each component is represented as a set of Gassusian primitives $\mathcal{G}_{i}$ parameterized with five attributes: 3D position $\mu_{i}\in\mathbb{R}^{3}$, opacity $\alpha_{i}\in\mathbb{R}$, rotation matrix $\mathbf{R}_{i}$ represented by axis angle $\mathbf{r}_{i}\in\mathbb{R}^{3}$, scale matrix $\mathbf{S}_{i}$ represented by diagonal vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ and rgb color $\mathbf{c}_{i}\in\mathbb{R}^{3}$. Inspired by existing works [22] that initialize Gaussian-based avatar by attaching them to SMPL parametric model. We initialize the 3D Gaussians of each component by attaching them to self-designed templates based on SMPL-X. As shown in Fig 3, we design templates for each component which can maximize the coverage of the region where the component may exist. The position of 3D Gaussian is initialized as the center point of faces on the template mesh. To enhance generation quality, all templates are subdivided to allow densified Gaussian primitives. We initialize the rotation of 3D Gaussians as the tangent frame of the faces consists of the normal vector of that face, the direction vector of one edge and their cross produce. The SMPL-X based initialization provides a shared geometry prior among subjects, facilitating reasonable geometry and easy animation.
@@
For the compatibility with feed-forward generation pipeline, previous methods usually project 3D Gaussians into 2D spaces(e.g. Multi-view image space [56, 54], UV space [72]). Different from existing gaussian-based methods, which represent the subject as an entity or adopting post-process for disentanglement, we directly model components by mapping their templates into a three-layer UV space combined with assigned semantic label designed by ourselves using Blender engine. The first layer represent the innermost human body, the second layer represent hair and shoes part and the third layer represent top and bottom clothes separately. All the attributes are encoded as local features, resulting in generated avatar with higher quality as stated in  [72]. We first separate the layered UV features plane into two parts in a channel-wise manner and decoded them with two light-weight shared decoders $\mathcal{D}_{g}$ and $\mathcal{D}_{t}$ separately. $\mathcal{D}_{g}$ decode the geometry-related attributes: position offset $\Delta\mu$ and opacity $\alpha$ of 3D Gaussians and the texture-related decoder $\mathcal{D}_{t}$ predict texture-related attributes, color $\mathbf{c}$ and covariance-related rotation $\Delta\mathbf{r}$ and scale $\Delta\mathbf{s}$. Given the decoded attribute maps, we can extract the attributes for each 3D Gaussian $\mathcal{G}_{i}$ via bilinear interpolation. The opacity value $\alpha_{i}$ and color value $\mathbf{c}_{i}$ are obtained directly and the other values are obtained via following formula based on their initial values:
$\mu_{i}$ $=\hat{\mu}_{i}+\Delta\mu_{i},$ (4)
$\mathbf{s}_{i}$ $=\hat{\mathbf{s}}_{i}\cdot\Delta\mathbf{s}_{i},$
$\mathbf{r}_{i}$ $=\hat{\mathbf{r}}_{i}\cdot\Delta\mathbf{r}_{i},$
@@
where $\hat{\mu}_{i}$, $\hat{\mathbf{s}}_{i}$ and $\hat{\mathbf{r}}_{i}$ are initial gaussian attribute values for position, scale and rotation separately as introduced before. $\Delta\mu_{i}$, $\Delta\mathbf{s}_{i}$ and $\Delta\mathbf{r}_{i}$ are predicted values extracted from attribute maps. After obtaining the attributes of each 3D Gaussian, we can obtain the generated digital avatars in a canonical space. By collecting the 3D Gaussians with the assigned semantic label, we can obtain representation of each component, and the clothed avatar can be represented as the composition of all the 3D Gaussians.
@@ 3.3 Deformation
Benefiting from the mentioned SMPL-X-based templates, our method enables deformation in body shapes and animation in novel poses including gestures and facial expressions. To support training with multiple subjects in various body shapes, we disentangle the body shape factor by defining all the templates in a canonical space with neutral body shapes. The neutral body shape avatar and its corresponding components can be transformed into targeted body shape space via the following warping process:
$\bar{\mu}=\mu+B_{s}(\mathbf{\beta},s,\mu),$ (5)
where $\bar{\mu}$ represents the position of 3D Gaussians in the targeted $\mathbf{\beta}$ body shape space and $B_{s}(\mathbf{\beta},s,\mu)$ are corresponding body shape related offsets extracted from the SMPL-X based templates via barycentric interpolation. We further add pose-dependent offsets $B_{p}(\mathbf{\theta},p,\mu)$ and facial expression offsets $B_{e}(\mathbf{\psi},e,\mu)$ in the same way to ensure accurate animation results.
The animation of the generated avatar from a canonical T-pose space into a targeted pose space can be regarded as transforming the 3D Gaussian attributes. During the animation process, the opacity $\alpha$ and color $c$ values of 3D Gaussians remain unchanged. Therefore, we only discuss the transformation of position $\mu$, rotation matrix $\mathbf{R}$ and scale matrix $\mathbf{S}$ in this section. Using the linear blend skinning technique, we transform the position $\mu$ of 3D Gaussians as:
$\mu^{\prime}=\sum_{i=1}^{n_{b}}w_{i}\mathbf{B}_{i}\bar{\mu},$ (6)
where $n_{b}$ represents the number of joints and $\mathbf{B}_{i}$ is the transformation matrix of the $i$-th joint. For 3D Gaussians in the innermost human body layer, the corresponding blend skinning weights $w$ are obtained directly from SMPL-X-based templates through barycentric interpolation as these regions usually undergo minimal topology changes. On the other hand, for 3D Gaussians representing the exterior layer, we extract the blending weights from a low-resolution volume containing precomputed fused skinning weights which is more robust for points far from the original template as stated by previous work [16]. The rotation matrix $\mathbf{R}$ is updated via $\mathbf{R}^{\prime}=\mathbf{T}_{1:3,1:3}\mathbf{R}$, where $\mathbf{T}=\sum_{i=1}^{n_{b}}w_{i}\mathbf{B}_{i}$. And the scale matrix $\mathbf{S}$ is recalculated in the targeted pose space.

### v2

@@ Abstract
Clothed avatar generation has wide applications in virtual and augmented reality, filmmaking, and more. While existing methods have made progress in creating animatable digital avatars, generating avatars with disentangled components (e.g., body, hair, and clothes) has long been a challenge. In this paper, we propose LayerAvatar, a novel feed-forward diffusion-based method capable of generating high-quality component-disentangled clothed avatars in seconds. We propose a layered UV feature plane representation, where components are distributed in different layers of the Gaussian-based UV feature plane with corresponding semantic labels. This representation can be effectively learned with current feed-forward generation pipelines, facilitating component disentanglement and enhancing details of generated avatars. Based on the well-designed representation, we train a single-stage diffusion model and introduce constrain terms to mitigate the severe occlusion issue of the innermost human body layer. Extensive experiments demonstrate the superior performances of our method in generating highly detailed and disentangled clothed avatars. In addition, we explore its applications in component transfer. The project page is available at https://olivia23333.github.io/LayerAvatar.
@@ 1 Introduction
Recent learning-based methods [3, 16, 84] mainly combine 3D representations [45, 71, 62, 30] with generation pipelines (e.g., 3D-aware GANs [49, 4, 5] and diffusion models [58, 53]) to create digital avatars. However, most of these methods often ignore the compositional nature of digital avatars and represent the human body, hair, and clothes as a whole, which limits their capabilities in digital avatar customization such as cloth transfer. Neural-ABC [7] and SMPLicit [14] provide parametric model with disentangled clothes and human body. However, modeling texture are leaved an un-explored problem. HumanLiff [24] proposes a layer-wise generation process that first generates clothed avatars in minimal clothes, then generates digital avatars wearing the next layer of clothing conditioned on the current layer. Nevertheless, the human body and clothes are not fully disentangled, which makes it difficult to extract the components of each layer, thus reducing the editing ability. In addition, some methods [70, 15, 66, 73, 19] follow the trend of DreamFusion [55] to achieve disentangled clothed avatar generation by learning each component of digital avatars separately through the prior knowledge of 2D diffusion models [58] in an optimization manner. These methods can generate clothed avatars with each component disentangled, however, they take hours to generate a single digital avatar, and the optimization time will increase linearly according to the number of components.
@@
In this paper, we propose LayerAvatar, a novel feed-forward diffusion-based method that achieves (1) component disentanglement, enabling seamless transfer of individual components such as clothes, hair, and shoes; (2) high-quality results, with generated avatars exhibiting intricate facial details, distinct fingers, and realistic cloth wrinkles; and (3) efficiency, requiring only seconds to generate a single avatar. We choose 3D Gaussians [30] as the underlying representation due to its high-quality rendering results for intricate details, and strong representation capability for diverse cloth types. However, naively representing the disentangled clothed avatar using 3D Gaussians is impractical due to its unstructured nature that is incompatible with most current feed-forward generation pipelines [5, 58]. Therefore, we introduce a Gaussian-based UV feature plane, in which 3D Gaussians are projected into a predefined 2D UV space shared among subjects. The attributes of each 3D Gaussian are encoded as local geometry and texture latent features, which can be obtained from the 2D feature plane via bilinear interpolation. Furthermore, to achieve full disentanglement of avatar components (hair, shoes, upper cloth) and higher generation quality, we represent avatar components in separate layers of the UV feature plane which provides neighboring components with distinctive features from different layers to facilitate decomposition.
@@
We evaluate LayerAvatar on multiple datasets [81, 9, 21], demonstrating its superior performance in generating disentangled avatars. We also explore its application in component transfer. In summary, our main contributions are:
  • We introduce LayerAvatar, a novel feed-forward clothed avatar generation pipeline with each component disentangled, enhancing the controllability of avatar generation.
  • We propose a layered UV feature plane representation that enhances generation quality and facilitates the disentanglement of each component.
  • Our method achieves outstanding generation results on multiple datasets and support downstream applications such as component transfer.
@@ 2 Related work
Figure 2: Method overview. LayerAvatar learns a feed-forward diffusion model to generate clothed avatars with each component disentangled. The clothed avatars are represented as layered UV feature plane where components are represented separately. After decoding the feature plane into attribute maps, we can extract 3D Gaussians from them through SMPL-X-based templates. Generated clothed avatars are then transformed into targeted pose space for further supervision. Reconstruction loss and constraint loss are both utilized to facilitate the disentanglement and handle the severe occlusion of human body layer.
@@
Compositional Avatar Representation. Instead of representing avatars [75, 80, 46] as a single entity, some methods represent clothed avatars as a combination of multiple submodules. COAP [44], Spams [51] and DANBO [65] consider human avatars as a composition of body parts, while EVA3D [22] and ENARF-GAN [50] follow this trend and utilize multiple neural networks to represent different body parts of the digital avatar, achieving more efficient and detailed generation results. Several methods [1, 78, 24] represent clothed avatars as separate layers to enable the expressiveness of various topologies. However, these methods ignore the disentanglement of human body and clothes, which makes each submodule less physically meaningful. Recently, some works [18, 54, 39, 77, 90] disentangle the human body and clothes by representing each component separately. During the rendering process, these components are combined through various compositional rendering techniques. Most of these methods are designed to optimize a single digital avatar. In this paper, we propose a novel layered UV feature plane representation that is compatible with feed-forward generation framework.
@@ 3 Method
We propose LayerAvatar, a feed-forward generative method for disentangled clothed avatar generation. The overview of our method is illustrated in Fig. 2. We provide a brief introduction for prior knowledge in Sec. 3.1. To achieve disentangled clothed avatar generation, we propose a novel layered UV feature plane representation (Sec. 3.2), which facilitates disentanglement and is compatible with current feed-forward generative pipelines. The clothed avatars are generated in canonical space and then deformed to targeted pose space via deformation module (Sec. 3.3). The training process is introduced in Sec. 3.4.
@@ 3.1 Preliminary
SMPL-X [52] is an expressive parametric human model that can produce naked human meshes $M(\mathbf{\beta},\mathbf{\theta},\mathbf{\psi})$ given shape parameter $\mathbf{\beta}$, pose parameter $\mathbf{\theta}$, and expression parameter $\mathbf{\psi}$. The producing process can be formulated as:
@@
3D Gaussians [30] is a primitive-based explicit representation that combines the strengths of both previous explicit and implicit representations. It consists of a set of learnable 3D Gaussian primitive $\mathcal{G}_{k}$ where each contains five attributes: position $\boldsymbol{\mu}$, scaling matrix $\mathbf{S}$, rotation matrix $\mathbf{R}$, opacity $\alpha$, and color $\mathbf{c}$. In practice, we employ diagonal vector $\mathbf{s}\in\mathbb{R}^{3}$ and axis-angle $\mathbf{r}\in\mathbb{R}^{3}$ to represent $\mathbf{S}$ and $\mathbf{R}$ respectively. 3D Gaussians are represented as ellipses in 3D space defined by their position $\mu$ and covariance matrix $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$. During the rendering process, these 3D Gaussians are projected to a 2D image plane where the pixel color $\mathbf{C}$ can be calculated as follows:
$\mathbf{C}=\sum_{i=1}^{N}\mathbf{c}_{i}\sigma_{i}\prod_{j=1}^{i-1}\left(1-\sigma_{j}\right),$ (2)
where $\mathbf{c}_{i}$ is the color of the $i$-th 3D Gaussian on the ray, and $\sigma_{i}$ is the blending weight calculated with the opacity $\alpha$.
@@ 3.2 Layered UV Feature Plane Representation
Previous methods mainly represent avatars as a single entity or depend on optimization-based schemes. As a result, the generated results often have difficulty with editing or are slow to create, typically taking hours to generate a single subject. To address these limitations, we propose a layered UV feature plane representation that can separate components of clothed avatars and is compatible with fast, feed-forward generation pipelines. We employ 3D Gaussians as the base representation for efficient rendering along with easy animation and editing.
@@
Each component is represented as a set of Gaussian primitives $\mathcal{G}_{i}$ parameterized with five attributes: 3D position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, opacity $\alpha_{i}\in\mathbb{R}$, rotation matrix $\mathbf{R}_{i}$ represented by axis angle $\mathbf{r}_{i}\in\mathbb{R}^{3}$, scale matrix $\mathbf{S}_{i}$ represented by diagonal vector $\mathbf{s}_{i}\in\mathbb{R}^{3}$ and rgb color $\mathbf{c}_{i}\in\mathbb{R}^{3}$. Inspired by existing works [25, 87] that initialize 3D Gaussians by attaching them to SMPL parametric model for geometry and animation prior. We initialize the 3D Gaussians of each component by attaching them to self-designed templates based on SMPL-X. For each component, we design a template that maximizes coverage of the region where the component may exist. To enhance generation quality, all templates are subdivided to support densified Gaussian primitives. Then, we initialize the positions of 3D Gaussians as the center points of faces on the densified template mesh. And the initial rotations of 3D Gaussians are set as the tangent frame of the faces, which consists of the normal vector of that face, the direction vector of one edge and their cross product.
@@
For compatibility with feed-forward generation pipelines, previous methods typically project 3D Gaussians into 2D space (e.g., multi-view image space [69, 67], UV space [87]). Unlike these methods, which represent the entire subject as a single entity and then employ a post-processing step for disentanglement, we directly model components separately by mapping their templates into a three-layer UV space, combined with semantic labels. The first layer represents the innermost part of the human body, the second layer represents the hair and shoes, and the third layer represents top and bottom clothes. Following  [87], all Gaussian attributes are stored as local UV features to enhance generation quality. The UV features plane is first split into two parts in a channel-wise manner and then decoded by two light-weight shared MLP decoders $\mathcal{D}_{g}$ and $\mathcal{D}_{t}$ separately. $\mathcal{D}_{g}$ decodes the geometry-related attributes: position offset $\Delta\boldsymbol{\mu}$ and opacity $\alpha$ of 3D Gaussians and $\mathcal{D}_{t}$ predicts texture-related attributes, color $\mathbf{c}$ and covariance-related rotation $\Delta\mathbf{r}$ and scale $\Delta\mathbf{s}$. Given the decoded attribute maps, we can extract the attributes for each 3D Gaussian $\mathcal{G}_{i}$ via bilinear interpolation. The opacity value $\alpha_{i}$ and color value $\mathbf{c}_{i}$ are obtained directly, and the other values are obtained via following formula based on their initial values:
$\boldsymbol{\mu}_{i}=\boldsymbol{\mu}^{0}_{i}+\Delta\boldsymbol{\mu}_{i},\quad\mathbf{s}_{i}=\mathbf{s}^{0}_{i}\cdot\Delta\mathbf{s}_{i},\quad\mathbf{r}_{i}=\mathbf{r}^{0}_{i}\cdot\Delta\mathbf{r}_{i},$ (4)
@@
where $\boldsymbol{\mu}^{0}_{i}$, $\mathbf{s}^{0}_{i}$, and $\mathbf{r}^{0}_{i}$ are initial position, scale, and rotation values for 3D Gaussians, respectively. $\Delta\boldsymbol{\mu}_{i}$, $\Delta\mathbf{s}_{i}$ and $\Delta\mathbf{r}_{i}$ are predicted residuals extracted from attribute maps. By collecting the attributes of 3D Gaussians with the same semantic label, we can obtain a canonical space representation of each component, and their composition forms the complete disentangled clothed digital avatar.
@@ 3.3 Deformation
Benefiting from the SMPL-X-based templates, our method supports deformation in body shapes and novel poses including gestures and facial expressions. To support training with multiple subjects in various body shapes, we disentangle the body shape factor by defining all the templates in a canonical space with neutral body shapes. The neutral body shape avatar and its corresponding components can be transformed into targeted body shape space via the following warping process:
$\bar{\boldsymbol{\mu}}=\boldsymbol{\mu}+B_{s}(\mathbf{\beta},s,\boldsymbol{\mu}),$ (5)
where $\bar{\boldsymbol{\mu}}$ represents the position of 3D Gaussians in the targeted $\mathbf{\beta}$ body shape space and $B_{s}(\mathbf{\beta},s,\boldsymbol{\mu})$ are corresponding body shape related offsets extracted from the SMPL-X based templates via barycentric interpolation. We further add pose-dependent offsets $B_{p}(\mathbf{\theta},p,\boldsymbol{\mu})$ and facial expression offsets $B_{e}(\mathbf{\psi},e,\boldsymbol{\mu})$ in the same way to ensure accurate animation results.
The animation of a generated avatar from the canonical T-pose to an arbitrary target pose can be regarded as transforming the 3D Gaussian attributes. During the animation process, the opacity $\alpha$ and color $\mathbf{c}$ of 3D Gaussians remain unchanged. Therefore, we only discuss the transformation of position $\boldsymbol{\mu}$, rotation matrix $\mathbf{R}$ and scale matrix $\mathbf{S}$ in this section. Using the LBS function, we can transform the position $\bar{\boldsymbol{\mu}}$ of 3D Gaussians as:
$\boldsymbol{\mu}^{\prime}=\sum_{i=1}^{n_{b}}w_{i}\mathbf{B}_{i}\bar{\boldsymbol{\mu}},$ (6)
where $n_{b}$ represents the number of joints and $\mathbf{B}_{i}$ is the transformation matrix of the $i$-th joint. For 3D Gaussians on the innermost human body layer, the corresponding blend skinning weights $w$ are obtained directly from SMPL-X-based templates through barycentric interpolation, as these regions usually undergo minimal topology changes. For 3D Gaussians representing the exterior components, we follow  [16] to extract the skinning weights from a pre-computed low-resolution volumetric field of fused skinning weights, which is more stable for points that deviate significantly from the original template. $\mathbf{T}=\sum_{i=1}^{n_{b}}w_{i}\mathbf{B}_{i}$ is the blended transformation matrix, and the rotation matrix $\mathbf{R}$ is updated via $\mathbf{R}^{\prime}=\mathbf{T}_{1:3,1:3}\mathbf{R}$, where $\mathbf{T}_{1:3,1:3}$ is the rotational part of $\mathbf{T}$. The scale matrix $\mathbf{S}$ is recalculated in the targeted pose space to fit deformed topology.